You want to buy your sister’s pickup truck that costs $4,800 all-in. Suppose she is
offering to finance it for you if you pay her each month with no money down, and 20%
APR for 2 years. What would be the monthly payment? What is the total amount paid
over the life of the loan? What is she getting in interest? She also offered to hold the car
for you for a year. How much would you have to save each month to be able to pay her
in full?
Answer #1

Monthly loan payment is calculated using PMT function in Excel :

rate = 20% / 12   (converting annual rate into monthly rate)

nper = 2*12 (2 year loan with 12 monthly payments each year)

pv = 4800 (amount financed)

PMT is calculated to be $244.30.

A1 for EPMT(20%/12,2*12,4800) D E F B C i ($244.30)

Total amount paid = monthly payment * number of payments.

Total amount paid = $244.30 * 2 * 12 = $5,863.20.

Interest = total amount paid - amount financed = $5,863.20 - $4,800 = $1,063.20.

Assuming that you can earn 20% interest on savings, the monthly saving to have $4,800 after 1 year is calculated using PMT function in Excel :

rate = 20%/12 (converting annual rate into monthly rate)

nper = 12 (12 monthly savings in 1 year)

pv = 0 (beginning savings are zero)

fv = 4800 (required savings after 1 year)

PMT is calculated to be $364.65.

A4 - fx =PMT(20%/12,12,0,4800) D E F B C ($364.65)!